# CoAR: Concept Injection into Autoregressive Models for Personalized Text-to-Image Generation

### v1

@@ Appendix C Additional Experiments
Figure 9: Qualitative results of subject-driven customization with CoAR.
Figure 10: Qualitative results of subject-driven customization with CoAR.
Additional Qualitative Results
@@
Figs.˜9, 10, 12, 11 and 13 presents additional qualitative results of subject-driven personalization and style personalization.
@@
Figs.˜9 and 10 demonstrate that given only a few reference images, CoAR successfully performs subject-driven generation across various tasks, including recontextualization (e.g., novel backgrounds), property modification (e.g., color and shape), and accessorization (e.g., adding glasses or outfits). The results demonstrate strong subject fidelity, adaptability to novel contexts, and fine-grained controllability.
Fig.˜12 presents examples of image generation conditioned on different artistic styles. Each row represents a different subject, while each column corresponds to a specific style. The generated images highlight the model’s ability to retain subject identity while adapting visual attributes such as texture, color palette, and rendering technique to match the target style.
@@
Fig.˜13 illustrates the effects of diverse contextual prompts on image generation for various subjects across different visual styles. Each row corresponds to a distinct subject, while each column varies the context.
@@
Figure 11: CoAR demonstrates effective style personalization conditioned on only a single reference image.
Figure 12: Style-conditioned image generation across subjects without contextual prompts.

### v2

@@ 8.1 Ablation Study on Dual Prior Preservation (DPP)
Figure 11: Qualitative ablation of token modalities across different insertion depths. Rows represent different token configurations, while columns correspond to the number of Transformer layers (1, 3, 9, 24) into which tokens are injected. Rows: “Only Text Token” suffers from identity loss (generic backpacks), while “Only Image Token” exhibits high fidelity but poor editability (e.g., failing to stylize in Col 4). Columns: Inserting tokens into all layers (Depth 24) leads to overfitting, preventing style changes, whereas the optimal depth (Depth 9) achieves the best balance between subject fidelity and textual control.
8.3 Joint Analysis of Modality and Insertion Depth
Figure 11 provides a comprehensive visualization of how different token modalities behave across varying insertion depths (injecting tokens into the first $N$ layers, where $N\in\{1,3,9,24\}$).
Impact of Modalities (Rows).
  • Only Text Tokens ($p_{[v]}$): As seen in the second row, relying solely on text tokens results in poor identity preservation across all depths. Although the text alignment is high (e.g., the “oil painting” style in Col 4 is correctly rendered), the generated backpacks are generic and lack the fine-grained details of the reference subject.
  • Only Image Tokens ($p_{I}$): The first row shows that while image tokens preserve visual details, they suffer from rigidity and poor instruction following, especially at deeper insertion layers.
  • Image and Text Tokens (Ours): The third row demonstrates that combining both modalities ensures robust identity preservation while remaining responsive to text prompts.
Impact of Insertion Depth (Columns).
  • Shallow Injection (Depth 1–3): In the early columns, the subject identity is not fully consolidated. For example, at Depth 1, the backpack’s texture and shape appear slightly inconsistent with the reference, indicating insufficient visual signal propagation.
  • Deep Injection (Depth 24): The last column (Depth 24) reveals the detrimental effect of over-injection. In the “Only Image Token” setting, the backpack fails to transform into the “oil painting” style and remains photorealistic. This indicates that injecting visual tokens into all layers creates an overly strong visual prior that overrides the textual style control, leading to overfitting.
  • Optimal Depth (Depth 9): Our chosen setting (Depth 9, Column 3) achieves the optimal sweet spot. The model successfully retains the specific identity of the backpack (unlike the text-only baseline) while seamlessly integrating it into the new background (unlike the depth-24 baseline), validating our decision to inject concepts only into the early-to-mid layers.
8.4 Additional Qualitative Results
@@
Figures˜14, 13, 15 and 16 presents additional qualitative results of subject-driven personalization and style personalization.
We provide further visual evidence comparing DCoAR with competing baselines in Figure˜12. Consistent with our main paper’s findings, shallow concept-injection methods (UniCTokens[UniCTokens], Yo’Chameleon[yoChameleon]) struggle to capture high-frequency details, often resulting in smoothed textures and lower visual fidelity. Conversely, while the adaptation-based method DreamBooth (utilizing the FLUX backbone) produces high-quality backgrounds, it exhibits significant semantic instability and identity drift; this is particularly evident in the second row, where the reference dog is erroneously rendered as a different breed. In contrast, DCoAR successfully preserves intricate subject details—such as specific fur patterns and material textures—while accurately adhering to complex background prompts, demonstrating the effectiveness of our Deep Concept Injection strategy.
Figure 12: Additional qualitative comparison of subject-driven personalization. We compare our DCoAR against two representative concept-injection methods (UniCTokens [UniCTokens], Yo’Chameleon [yoChameleon]) and a state-of-the-art diffusion-based fine-tuning approach (DreamBooth [ruiz2023dreambooth] with FLUX).
@@
Figures˜14 and 13 demonstrate that given only a few reference images, DCoAR successfully performs subject-driven generation across various tasks, including recontextualization (e.g., novel backgrounds), property modification (e.g., color and shape), and accessorization (e.g., adding glasses or outfits). The results demonstrate strong subject fidelity, adaptability to novel contexts, and fine-grained controllability.
Figure 13: Qualitative results of subject-driven customization with DCoAR.
@@
Figure˜16 illustrates the effects of diverse contextual prompts on image generation for various subjects across different visual styles. Each row corresponds to a distinct subject, while each column varies the context.
9 Limitations
Despite the superior performance of DCoAR in personalized generation, we acknowledge the primary limitations that warrant future investigation: Sensitivity to Insertion Depth and Overfitting. Our framework relies on a carefully tuned Layer-wise Multimodal Context Learning (LMCL) strategy. As evidenced by our ablation studies, the model is sensitive to the depth of token injection. While shallow injection fails to capture identity details, injecting tokens into overly deep layers imposes excessive visual constraints. This leads to overfitting, where the model becomes rigid and prioritizes pixel-level adherence to the reference image over the semantic control of the text prompt, hindering effective re-contextualization and stylization. Developing an adaptive mechanism to automatically determine the optimal insertion depth for different subjects remains an open problem.
Figure 14: Qualitative results of subject-driven customization with DCoAR.
@@ 9 Limitations
Figure 15: DCoAR demonstrates effective style personalization conditioned on only a single reference image.